ABC’s income has increased from P 20,000 to P 30,000 per month. Because of this, his consumption for fruits has increased by 20%. Compute the income elasticity and determine the type of good.


1
Expert's answer
2021-09-15T16:17:20-0400

By the definition of the income elasticity of demand we have:


"E_Y=\dfrac{\%\Delta Q}{\%\Delta Y}=\dfrac{\%\Delta Q}{\dfrac{Y_2-Y_1}{Y_1}\cdot100\%},""E_Y=\dfrac{20\%}{\dfrac{30000-20000}{20000}\cdot100\%}=0.4"

As we can see from calculations, the income elasticity of demand is positive and lies between 0 and +1, therefore it is a normal good (necessity).